# The ML Supply Chain in the Era of Software 2.0: Lessons Learned from Hugging Face

### v1

@@ 7.1. License compliance in the ML supply chain
In Section 5, we showed that the structure of the ML supply chain consists primarily of relatively short component chains. This presents a radically different landscape from the traditional software supply chain, in which a single piece of software can depend on potentially hundreds or even thousands of components. However, just because fewer dependencies are declared does not mean that ensuring license compliance is any easier in this ecosystem. In fact, as we observed, licensing information is often missing (Section 6) and there can be licensing differences between parent and child models in a model’s lineage (Section 6.2). Compounded with our observation that datasets may also have unclear licensing (Section 6.1) and may go undeclared by models (Section 5.2), the state of documentation on Hugging Face appears to be counterproductive to license compliance tasks.
@@
It also becomes crucial to educate model owners and users on the terms associated with licenses. For example, we observed that the simple naming requirements imposed by the Llama3 license were followed in only 14.2% of cases (Section 6.4.2). This suggests either extensive noncompliance or, more likely, an ignorance of the terms being violated. To concretely identify the cause and provide potential mitigations, future work should investigate developers’ views and understanding of licenses in the ML supply chain.
@@
It also remains unclear, particularly with the relationship between datasets and models, how exactly license terms, such as GPL’s copyleft, should be applied. To what extent should a model be considered a derivative work of the dataset(s) on which it was trained? More concretely, if a model trainer trains a model on a dataset governed by a copyleft license and then distributes the model, but not the dataset, has the trainer violated the dataset license? These questions and many others will likely need to be resolved by legal teams, regulators, and potentially the courts.
@@
Lastly, in Section 6.4.1, we also observed instances of ambiguity introduced through multi-licensing. Determining the intended relationship among the multiple licenses assigned by the model/dataset owner can be a difficult task, and different pitfalls come with the assumption of an OR versus an AND relationship. Model hubs should include standardized ways for model/dataset owners to express these relationships and educate users on why they might want to use a multi-licensing scheme. It is left to future work to explore why some developers are presently releasing their models/datasets under multiple licenses.
7.2. Potential difficulties in using model cards as ML/AIBOM
@@
Amidst the growing push for better transparency and security in software through Software Bills of Materials (SBOMs) (Xia et al., 2023a; Stalnaker et al., 2024), calls have been made for similar BOMs for ML components (Stalnaker et al., 2024; Xia et al., 2023a; Barclay et al., 2019, 2022). These documents serve as inventories of all components within a piece of software and can document a variety of the software’s traits, including dependencies, licensing, and security. Distinct ML/AI Bills of Materials (ML/AIBOMs) may be necessary to address such components’ different inputs, security concerns (such as model poisoning), and ethical considerations (Xia et al., 2023a, b; Bi et al., 2024; Lu et al., 2024). The SPDX (spdx, 2025) Working Group, operating under the auspices of the Linux Foundation, is developing guidelines and a proposed standard for such an ML/AIBOMs (Bennet et al., 2024) that will complement the SPDX ISO Standard ISO/IEC 5962:2021, which describes the use of SBOMs to document the components used in creating a software system (ISO/IEC 5962:2021, 2021). CycloneDX (cyclonedx, 2025a) is also working on support for ML/AIBOMs (cyclonedx, 2025b).
@@
Prior work has also suggested that model and data cards can serve as an ML/AIBOM (Stalnaker et al., 2024), or the information they provide can assist with their creation (Xia et al., 2023b; Pepe et al., 2024), but our work suggests that these tools are not yet robust enough to serve this purpose. As we highlight in Section 4, the information provided by model cards is often missing key elements, including datasets that the models were trained on. This immediately obviates many of the benefits of ML/AIBOMs, including understanding licensing obligations that might be associated with that dataset, being aware of potential model poisoning attacks, and providing the ability to select models trained on ethically sourced data. Additionally, the cases where model cards were locked behind Terms of Service agreements or other restrictions (Section 4.1) could further limit their usefulness as ML/AIBOMs to consumers. Our work provides additional evidence that, in practice, model cards contain little actionable information, making them difficult to use as ML/AIBOMs (Bhat et al., 2023).
@@ 7.3. The need for follow up surveys and interviews
Our study raises many questions concerning what motivates developers when they make decisions concerning documentation, model reuse, and licensing. Conversations with Hugging Face developers through surveys and interviews can help to bridge this gap. There are numerous key questions such work should consider. What factors influence developers when choosing a license for their dataset/model (Section 6.1)? Why do developers so commonly opt to use an “other” license (Section 6.1)? What motivates developers to multi-license a model/dataset and what is the intention behind the license interaction (Section 6.4.1)? Why do so few developers fine-tuning the llama3 model comply with the naming requirements imposed by the llama3 license (Section 6.4.2)? What is the primary motivation for developers to write documentation for their models? For each of the various stakeholders in the space (model owners, fine-tuners, and users), which metadata fields do they perceive as most important? Why are some models listed as datasets (Section 4.3.4)? Answers to these questions will inform how future tools, practices, techniques, and even educational materials should be developed to have maximal impact on the emerging ML community.
@@ 7.4. The addition of fields for recording model relationship information
As the relationships between ML models become more complex, it becomes increasingly important to keep a record of those relationships (e.g., fine-tuned, quantized, ensemble, student/teacher, etc.). This information is crucial not only for transparency and understanding how the model was created, but also for license compliance. Just as in traditional software, licenses may have terms that govern the manner in which the component may be (re)used and distributed. For example, many generative models disallow using generated content to train a derivative model (openai-terms, 2023). As noted in Section 4.3.5, Hugging Face provides no standardized way to express relationship information. This inability to specify relationships may have resulted in some model owners declaring models used in student-teacher relationships as datasets (Section 4.3.4). We suggest that model hubs provide a standardized, machine-readable way of conveying this type of information. Just as Hugging Face currently provides a standard list of architectures, it could also provide a curated list of relationship types.
@@ 7.5. The need to use unique identifiers
While unique identifiers already exist on Hugging Face, they are not currently being used by model/dataset owners. This can result in problems, particularly when trying to trace model/dataset provenance. Model hubs should encourage their users to use these existing unique identifiers, perhaps by including tools that facilitate the creation of model cards or the provision of metadata. For example, a developer could supply a colloquial name for a model hosted on the platform, such as the familiar owner/model convention, and the tool could automatically conduct a search and, where possible, map the provided name to the unique ID. This would immediately alert the developer of typos, remove ambiguities, and also prevent the name drift issues we discussed in Sections 4.3.2 and 4.3.3. Furthermore, such a system could potentially allow for automatic updates (or request manual developer updates) when the information for downstream elements changes, thereby facilitating an ecosystem which is more supportive of supply chain management tasks and lifting the burden from downstream developers.
@@ 7.6. Differences between the ML supply chain and traditional software supply chain
Our findings highlight a notable difference between the ML supply chain and the traditional software supply chain. In many software ecosystems with dependency management, dependency information for components in the software supply chain is organized in manifest files such as pom.xml, package.json, or requirements.txt. These files can then be used to generate other dependency tracking documents like SBOMs and have also been used by forges like GitHub to build dependency graphs (dep, [n. d.]). Notably, however, these manifest files have utility beyond dependency tracking. They are necessary for setting up a fresh installation of the software – that is, the list of dependencies must be downloaded in order for the software to function properly. This is not the case with the information supplied by developers in model cards. Because erroneous mappings to datasets and base models do not leave the model in question unusable as a practical matter, it is therefore easier for typos and other mistakes to go unnoticed, and thus uncorrected. This disconnect between correctness and functionality necessitates tooling that has similar functionality and richness to dependency management tools for traditional software, as it leaves ample opportunity for errors without comparable means of checking for and correcting them.
@@ 7.7. Quality control for model hubs
Our analysis of the model documentation available on Hugging Face shows that the information that developers provide is often incomplete, malformed, or only available in human-readable text. We observed some documentation issues, such as the presence of duplicates in lists of datasets and base models (Section 4.3.5) or the declaration of a model as its own base (Section 4.3.3), that could be easily corrected by automated tools like linters. Additionally, an automated tool could also provide standardization for data fields to facilitate metadata analysis. For example, the variants for how one can specify that there are no datasets or base models ([], “”, and None) should be standardized to a single type. We are not suggesting that model hubs should be responsible for curating the metadata of all models on their platforms, but where automated checks can efficiently be put into place, they should be employed.
Potential solutions could closely resemble approaches to automatically check the quality of bug reports. Similarly to how tools can check for important components of a bug report (e.g.,, expected behavior, steps to reproduce, etc.(Chaparro et al., 2017), in theory, tools may be able to check for missing information in model cards. Some work is already emerging in this area, such as DocML (Bhat et al., 2023) and AIMML (Tsay et al., 2020), however, more work needs to be done to understand what suggestions should be made to model owners, how to handle complex licensing scenarios, and how to best implement tools that assess ML documentation.
@@ 7.8. Potential pitfalls when mining Hugging Face
The numerous documentation challenges that we discuss in Section 4, demonstrate that researchers relying on Hugging Face mining to collect information on the ML supply chain must proceed thoughtfully, with caution and care to avoid potential pitfalls. Unlike source code forges like GitHub, it seems that some model owners may be uploading and publishing models as a by-product of their workflows. As such, we speculate that some models are made publicly available not because the owner intends them to be reused by the community, but rather because the owner is using the resources of Hugging Face, Google Colab, or some other cloud provider. (This might explain why we observed so many models with the same names. See Section 4.3.2.) More research is required to understand how such tools might populate the supply chain with such “intermediate” models, or those that are otherwise not intended for public consumption. To mitigate the risk such models pose, future studies should be mindful of the models they include in their analysis, erring toward those that are more popular. However, determining popularity may not be as easy as sorting by downloads or likes. As shown in Table 7, the most liked “dataset” on Hugging Face is in reality a repository curating ChatGPT prompts. Future studies will be left to explore the prevalence of curated repositories hosted as datasets and the exact nature of potentially ad hoc developer workflows.

### v2

@@ 10.1. License compliance in the ML supply chain
In Section 6, we showed that the structure of the ML supply chain on Hugging Face consists primarily of relatively short component chains. This presents a radically different landscape from the traditional software supply chain, in which a single piece of software can depend on potentially hundreds or even thousands of components. However, just because fewer dependencies are declared does not mean that ensuring license compliance is any easier in this ecosystem. In fact, as we observed, licensing information is often missing (Section 7) and there can be licensing differences between parent and child models in a model’s lineage (Section 7.3). Compounded with our observation that datasets may also have unclear licensing (Section 7.1) and may go undeclared by models (Section 6.2), the state of documentation on Hugging Face appears to be counterproductive to successful license compliance.
@@
It also becomes crucial to educate model owners and users on the terms associated with licenses. For example, we observed that the simple naming requirements imposed by the Llama3 license were followed in only 14.2% of cases (Section 7.5.2). This suggests either extensive deliberate noncompliance or, more likely, ignorance of the relevant license terms. To concretely identify the cause and provide potential mitigations, future work should investigate developers’ views and understanding of licenses in the ML supply chain.
@@
It also remains unclear, particularly with the relationship between datasets and models, how exactly license terms, such as GPL’s copyleft, should be applied (Hardy, [n. d.]). To what extent should a model be considered a derivative work of the dataset(s) on which it was trained? More concretely, if a model trainer trains a model on a dataset governed by a copyleft license and then distributes the model, but not the dataset, has the trainer violated the dataset license? More fundamentally, to what extent are ML datasets protected by copyright law in various jurisdictions? In other words, is a particular selection and arrangement of data protected by copyright, or does copyright attach only at the level of individual data, if at all? These questions and many others will likely need to be resolved by legal teams, regulators, and possibly the courts.
We observed challenges presented by licensing datasets that contain data points under different licenses (Sections 9.3.3 and 9.3.4). While these challenges are not new (Longpre et al., 2023; Lee et al., 2023b), future research and the Hugging Face development team should investigate more intuitive ways of conveying that the licensing of a dataset depends on the specific portions used rather than allowing users to obscure this information by the use of an “Other” label, instead reserving the “Other” label for custom or non-standard licenses. Relatedly, dataset and model repositories should encourage model/dataset uploaders to select from already existing licenses rather than using “Other” in order to avoid rampant license proliferation. If custom license text is required, it should be within a LICENSE file so that it is easy for users to locate and access.
@@
Lastly, in Sections 7.5.1, 9.3.5 and 8.4.1, we also observed instances of ambiguity introduced through multi-licensing. Determining the intended relationship among the multiple licenses assigned by the model/dataset owner can be a difficult task, and different pitfalls come with the assumption of an OR versus an AND relationship. Model hubs should include standardized ways for model/dataset owners to express these relationships and educate users on why they might want to use a multi-licensing scheme.
10.2. Dataset aggregates
As discussed in Sections 9.3.3 and 9.3.4, some datasets on Hugging Face were aggregates or composites of other datasets. In these situations, the uploader would leave license identification to the end user, saying something to the effect of “Look at the licenses for the different subsets of this dataset.” In addition to raising potential compliance questions, the lack of standardization in the way that this aggregate nature was conveyed potentially presents challenges. We observed model uploaders using the “Unknown” and “Other” license options to refer to this composite nature. Other uploaders opted to leave their dataset unlicensed in the tags and elaborated in the body of the dataset card that the terms of the individual licenses needed to be complied with. Still others opted to declare multiple licenses in their tags, not to indicate an AND or an OR relationship but rather to indicate that different licenses applied to different subparts of the dataset (Section 9.3.5). Without an additional field to allow owners to describe this type of licensing situation more precisely, this structure is likely to contribute to developers’ difficulty with understanding licensing tasks, a difficulty that previous work has explored (Almeida et al., 2019).
The purpose of these dataset aggregates would seem to be assembling large quantities of training data for machine learning models and making them easily accessible. Additional problems arise if the licenses contained within an aggregate conflict. For example, nguha/legalbench is composed of datasets that are under two different copyleft CC licenses (Attribution-ShareAlike 4.0 International and Attribution-NonCommercial-ShareAlike 4.0 International). It is not possible to meet the terms of both licenses, meaning that this dataset composite and models trained on this dataset risk being deemed noncompliant. This also poses risks for any end user, who might not have the motivation or the knowledge to determine compliance, raising nonfunctional ethical considerations that are deserving of attention by researchers.
10.3. Potential difficulties in using model cards as ML/AIBOMs
@@
Amidst the growing push for better transparency and security in software through Software Bills of Materials (SBOMs) (Xia et al., 2023a; Stalnaker et al., 2024), calls have been made for similar BOMs for ML components (Stalnaker et al., 2024; Xia et al., 2023a; Barclay et al., 2019, 2022). These documents serve as inventories of all components within a piece of software and can document a variety of the software’s traits, including dependencies, licensing, and security. Distinct ML/AI Bills of Materials (ML/AIBOMs) may be necessary to address such components’ different inputs, security concerns (such as model poisoning), and ethical considerations (Xia et al., 2023a, b; Bi et al., 2024; Lu et al., 2024). The SPDX (spdx, 2025) Working Group, operating under the auspices of the Linux Foundation, is developing guidelines and a proposed standard for ML/AIBOMs (Bennet et al., 2024) that will complement the SPDX ISO Standard ISO/IEC 5962:2021, which describes the use of SBOMs to document the components used in creating a software system (ISO/IEC 5962:2021, 2021). CycloneDX (cyclonedx, 2025a) is also working on support for ML/AIBOMs (cyclonedx, 2025b).
@@
Prior work has also suggested that model and data cards can serve as an ML/AIBOM (Stalnaker et al., 2024), or that the information they provide can assist with their creation (Xia et al., 2023b; Pepe et al., 2024), but our work suggests that these tools are not yet robust enough to serve this purpose. As we highlight in Section 5, the information provided by model cards is often missing key elements, including datasets that the models were trained on. This immediately obviates many of the benefits of ML/AIBOMs, including understanding licensing obligations that might be associated with that dataset, being aware of potential model poisoning attacks, and providing the ability to select models trained on ethically sourced data. Additionally, instances in which model cards are locked behind Terms of Service agreements or other restrictions (Section 5.1) could further limit their usefulness as ML/AIBOMs to consumers. Our work provides additional evidence that, in practice, model cards contain little actionable information, making them difficult to use as ML/AIBOMs (Bhat et al., 2023).
@@ 10.4. The need for follow up surveys and interviews
Our study raises many questions concerning what motivates developers when they make decisions concerning documentation, model reuse, and licensing. Conversations with Hugging Face developers through surveys and interviews can help to bridge this gap. There are numerous key questions such work should consider. What factors influence developers when choosing a license for their dataset/model and what do they believe the scope of that license to be (Section 7.1)? Why do developers so commonly opt to use an “Other” license (Section 7.1)? What motivates developers to multi-license a model/dataset and what is the intention behind the license interaction (Section 7.5.1)? Why do so few developers fine-tuning the llama3 model comply with the naming requirements imposed by the llama3 license (Section 7.5.2)? What is the primary motivation for developers to write documentation for their models? For each of the various stakeholders in the space (model owners, fine-tuners, and users), which metadata fields do they perceive as most important? Why are some models listed as datasets (Section 5.6)? Answers to these questions will inform how future tools, practices, techniques, and even educational materials should be developed to have maximal impact on the emerging ML community.
@@ 10.5. The addition of fields for recording model relationship information
As the relationships between ML models become more complex, it becomes increasingly important to keep a record of those relationships (fine-tuned, quantized, ensemble, distillation, etc.). This information is crucial not only for transparency and understanding how the model was created but also for license compliance. Just as in traditional software, licenses may have terms that govern the manner in which the component may be (re)used and distributed. For example, many generative models disallow using generated content to train a derivative model (openai-terms, 2023). As noted in Section 5.7, Hugging Face provides no standardized way to express relationship information. This inability to specify relationships may have resulted in some model owners declaring models used in student-teacher relationships as datasets (Section 5.6). We suggest that model hubs provide a standardized, machine-readable way of conveying this type of information. Just as Hugging Face currently provides a standard list of architectures, it could also provide a curated list of relationship types.
@@ 10.6. The need to use unique identifiers
While unique identifiers already exist on Hugging Face, they are not currently being used by model/dataset owners. This can result in problems, particularly when trying to trace model/dataset provenance. Model hubs should encourage their users to use these existing unique identifiers, perhaps by including tools that facilitate the creation of model cards or the provision of metadata. For example, a developer could supply a colloquial name for a model hosted on the platform, such as the familiar owner/model convention, and the tool could automatically conduct a search and, where possible, map the provided name to the unique ID. This would immediately alert the developer to typos, remove ambiguities, and prevent the name drift issues we discussed in Sections 5.4 and 5.5. Furthermore, such a system could potentially allow for automatic updates (or request manual developer updates) when the information for downstream elements changes, thereby facilitating an ecosystem which is more supportive of supply chain management tasks and lifting the burden from downstream developers.
@@ 10.7. Differences between the ML supply chain and traditional software supply chain
Our findings highlight a notable difference between the ML supply chain and the traditional software supply chain. In many software ecosystems with dependency management, dependency information for components in the software supply chain is organized in manifest files such as pom.xml, package.json, or requirements.txt. These files can then be used to generate other dependency tracking documents like SBOMs and have also been used by forges like GitHub to build dependency graphs (dep, [n. d.]). Notably, however, these manifest files have utility beyond dependency tracking. They are necessary for setting up a fresh installation of the software—that is, the list of dependencies that must be downloaded in order for the software to function properly. This is not the case with the information supplied by developers in model cards. Because erroneous mappings to datasets and base models do not leave the model in question unusable as a practical matter, it is therefore easier for typos and other mistakes to go unnoticed and thus uncorrected. This disconnect between correctness and functionality necessitates tooling that has similar functionality and richness to dependency management tools for traditional software, as it leaves ample opportunity for errors without comparable means of checking for and correcting them.
@@ 10.8. Quality control for model hubs
Our analysis of the model documentation available on Hugging Face shows that the information that developers provide is often incomplete, malformed, or available only in human-readable text. We observed some documentation issues, such as the presence of duplicates in lists of datasets and base models (Section 5.7) or the declaration of a model as its own base (Section 5.5) that could be easily corrected by automated tools like linters. Additionally, an automated tool could also provide standardization for data fields to facilitate metadata analysis. For example, the variants for how one can specify that there are no datasets or base models ([], “”, and None) should be standardized to a single type. We are not suggesting that model hubs should be responsible for curating the metadata of all models on their platforms, but where automated checks can efficiently be put into place and this is easy to implement, they should be employed.
Potential solutions could closely resemble approaches to automatically check the quality of bug reports. Similarly to how tools can check for important components of a bug report (e.g., expected behavior, steps to reproduce) (Chaparro et al., 2017), in theory, tools may be able to check for missing information in model cards. Some work is already emerging in this area, such as DocML (Bhat et al., 2023) and AIMML (Tsay et al., 2020); however, more work needs to be done to understand what suggestions should be made to model owners, how to handle complex licensing scenarios, and how to best implement tools that assess ML documentation.
Since license metadata for models and datasets is entered manually, there is the possibility that typos are introduced. We encountered instances where the standard SPDX identifier cc-by-sa-4.0 was mistakenly entered as cc-by-sa 4.0 (i.e., missing the final hyphen). While not a major issue, typos like these necessitate that tools perform data cleaning in addition to any work related to license compliance. Again, it would be preferable if Hugging Face provided a basic linter/validator for model/dataset cards and provided users with a drop-down/auto-complete from which to select popular licensing options, if this would be easy to implement.
@@ 10.9. Potential pitfalls when mining Hugging Face
The numerous documentation challenges that we discuss in Section 5 demonstrate that researchers relying on Hugging Face mining to collect information on the ML supply chain must proceed thoughtfully, with caution and care, to avoid potential pitfalls. Unlike source code forges like GitHub, it seems that some model owners may be uploading and publishing models as a by-product of their workflows (Face, 2025d, b). As such, we speculate that some models are made publicly available not because the owner intends them to be reused by the community but rather because the owner is using the resources of Hugging Face, Google Colab, or some other cloud provider (Boudier and Schmid, [n. d.]). This might explain why we observed so many models with the same names, such as my_awesome_model which is the example name used in the documentation for uploading a PyTorch model (Face, 2025d). (See Section 5.4.) More research is required to understand how such tools might populate the supply chain with such “intermediate” models or those that are otherwise not intended for public consumption. To mitigate the risk such models pose, future studies should be mindful of the models they include in their analysis, erring toward those that are more popular. However, we caution researchers to be thoughtful in how they determine model popularity. In the context of the supply chain, we believe that downloads is the more appropriate metric, as opposed to likes. As shown in Table 7, the most liked “dataset” on Hugging Face is in reality a repository curating ChatGPT system prompts (e.g., “Imagine you are an experience Ethereum developer…”). Additionally, the repository for lllyasviel/sd_control_collection (with 1,574 likes, but 0 downloads) is actually a collection of various models assembled “for users to download flexibly.” Future studies will be left to explore the prevalence of curated repositories hosted as models/datasets and the exact nature of potentially ad hoc developer workflows.
10.10. Recommendations for Stakeholders
Here, we offer concise and actionable recommendations for various stakeholders involved in the ML supply chain based on our results.
10.10.1. Model trainers and dataset curators
  1. (1)
    When licensing models and datasets, use tags, the CardData attribute, and/or a LICENSE file rather than including licensing information in the README or via an external source.
  2. (2)
    When releasing a model or dataset under a multi-licensing scheme, clearly indicate how the licenses relate.
  3. (3)
    When choosing a license for models and datasets, where possible use common, pre-existing licenses, rather than drafting a new license or modifying the terms of an existing one.
  4. (4)
    When there is no choice but to draft a new license, make sure that the full license text is made available in a LICENSE file available on the model/dataset repository. Also ensure that licensing terms are unambiguous and comprehensive (e.g., more than “This model is open source”).
  5. (5)
    When using models and datasets, particularly for larger or commercial projects, verify that the declared license is consistent and compliant with other upstream sources.
10.10.2. Model and dataset sharing platforms
  1. (1)
    Provide intuitive ways for dataset curators to specify the licenses of their work, particularly in the case of dataset aggregates. Simply using the “Other” or “Unknown” license options or declaring multiple licenses without clarification seems insufficient.
  2. (2)
    When designing documentation tools for model trainers and dataset curators, offer user-friendly options, such as multi-select widgets ( with an “Other” option) and smart auto-complete text fields, instead of simple plain text fields. This is particularly the case for for fields that have a finite set of recognizable values, such as licenses, pipelines, etc.
  3. (3)
    Perform basic validation of model/dataset metadata, including checking for obvious typos, needless repetitions, or improper uses of fields.
  4. (4)
    Automatically include unique identifiers for models and datasets listed as dependencies when possible. Human readable names for both models and datasets can change over time, but humans are also not well suited for working with the random strings of characters often associated with unique identifiers. When models/datasets are referenced in the metadata of other models/datasets, they should also be accompanied with a unique identifier that anchors the dependency, even if the human readable name should change. For Hugging Face, this might mean automatically adding the internal model ID as part of the base_model information in the CardData attribute.
  5. (5)
    Offer standardized, precise ways for model trainers to specify the dependency relationships between models (fine-tune, distillation, quantization, etc.), as these may have implications for various supply chain management activities, including licensing compliance.
  6. (6)
    For platforms that host both models and datasets, allow model trainers to specify where datasets are located with an external link if they are not available directly on the platform in order to make various provenance tasks significantly easier.
10.10.3. SE and Law Researchers
  1. (1)
    Be aware and cautious of the limitations and quirks of mining data from Hugging Face. For example, the readily available “downloads” metadata field actually only reports downloads that have occurred within the last 30 days. Total downloads can be accessed on a per query basis by providing additional arguments through the API, but it is not possible (as far as the authors can tell) to have the API return a list of models ordered by total downloads.
  2. (2)
    Choose proxy metrics carefully. As described in Section 10.9, when using model/dataset popularity/engagement as a proxy for quality or importance, be mindful that like count (similar to GitHub’s stars) may not be as reliable as the download count.
  3. (3)
    Understand that dependency information in the ML supply chain is more difficult to validate than dependency information in the traditional supply chain as described in Section 10.7.
  4. (4)
    Develop updated guidance and compliance matrices that can support stakeholders in the ML ecosystem as they work with various different licenses in a novel context.
  5. (5)
    Investigate developers’ views and understanding of licenses in the ML supply chain through surveys and interviews.
  6. (6)
    Know that not all models on Hugging Face were uploaded with the intent to publish. Instead, some may be incidental uploads that are a by-product of developer workflows through which they access cloud computing resources.